# Zonal vs. Nodal Pricing: An Analysis of Different Pricing Rules in the German Day-Ahead Market

### v1

@@ 2.2.1 Zonal Market Clearing
The key idea of zonal pricing is to aggregate nodes into zones and only consider transmission flows across zones. This simplifies the constraint set $\Psi$ significantly, enhances computational scalability, and allows for portfolio bidding of larger market participants. In the European market, two approaches have been implemented: the available transfer capacity (ATC) model and the flow-based market coupling (FBMC) methodology (NEMO Committee,, 2019). Both approaches require a refined capacity calculation process before the allocation to assess the available cross-zonal flow capacities. Furthermore, the FBMC approach necessitates access to a power transfer distribution factor (PTDF) matrix that indicates how zone net position changes affect cross-zonal flows.
As these data were not included in the BZR data release, we rely on a different approach to compute a zonal allocation. Specifically, to determine the aggregate characteristics of a cross-zonal connection, we sum up the thermal limits and use the mean susceptance over all relevant cross-zonal lines. Specifically, with a set of zones $Z$ and $B_{zz^{\prime}}=\overline{B}_{nm:n\in z,m\in z^{\prime}}$ as the mean cross-zonal susceptance, the zonal constraint set is defined as follows.
$\Psi^{DC}_{Zonal}=\{(x,y):\sum_{s\in S}y_{s,zt}-\sum_{b\in B}x_{b,zt}=\sum_{z^{\prime}\in Z}-B_{zz^{\prime}}(\phi_{z}-\phi_{z}^{\prime})\;\forall z\in Z,t\in T\}$ (5)
In essence, (5) is a natural extension of (4) where a set of aggregate zones is considered instead of nodes. Cross-zonal flows are feasible if they satisfy $\Psi^{DC}_{Zonal}$.
@@ 2.2.2 Redispatch
This work simulates a cost-based redispatch in line with real-world practices. As redispatch should be cost-minimal, the fundamental idea of our approach is to find a modified allocation that satisfies the constraint set of $\Psi^{DC}$ and, at the same time, minimizes the required compensations to sellers for their changes in generation. Demand is excluded from redispatch, as it is current practice in Germany. More formally, with $(x^{Zonal},y^{Zonal})$ as the outcome of the zonal allocation, we solve the following problem:
@@
subject to $x=x^{Zonal}$
$x,y\in\Psi^{DC}$
The objective of the optimization problem described in (6) is to adjust the zonal allocation $(x^{Zonal},y^{Zonal})$ in a manner that yields a DCOPF-feasible solution $(x,y)$ at minimal costs. This adjustment involves compensating sellers for any changes in their dispatched quantities during redispatch. For each seller $s$, this compensation is determined by the difference between the costs they incur under the initial zonal allocation $c_{s}(y_{s}^{Zonal})$ and the costs under the final DCOPF-feasible allocation $c_{s}(y_{s})$. The objective of (6) is to minimize the total costs associated with these adjustments. As part of our experiments we also report results for redispatch that minimizes volumes instead of costs, i.e., the redispatch objective function changes to $\min_{x,y}\;\sum_{s\in S}\lvert y_{s}-y_{s}^{Zonal}\rvert$.
@@ 2.3 Pricing on Non-Convex Markets
Consequently, market operators have to resort to heuristic pricing rules that aim at approximating Walrasian equilibrium prices. Many such pricing rules have been proposed (Liberopoulos & Andrianesis,, 2016), each compromising on the properties of Walrasian equilibria differently. In this work, we focus on Convex Hull (CH) pricing (Hogan & Ring,, 2003; Gribik et al.,, 2007) and Integer Programming (IP) pricing (O’Neill et al.,, 2005) as two established pricing rules in practice, as well as on the Join pricing rule (Ahunbay et al.,, 2022) as a novel multi-objective approach to pricing. We also consider the Euphemia pricing rule as it is applied in European electricity markets.
@@ 2.3.2 IP Pricing
[…]
By convexity of $v_{b}(x_{b})$, a buyer’s LLOCs exactly equal their GLOCs. Besides computational tractability, IP prices accurately reflect the marginal value of transmission capacity (Yang et al.,, 2019). Specifically, price differences among nodes arise solely when the network experiences congestion. It has been shown that this property immediately corresponds minimizing LLOCs (Ahunbay et al.,, 2022), making them a significant indicator of good congestion signals.
@@ 2.3.3 Join Pricing
[…]
Bids that require MWPs are also known as paradoxically accepted bids. MWPs can be regarded as another subset of GLOCs, referring only to participants’ incentives to not participate in the market at all. However, both CH and IP pricing can experience high levels of MWPs. Recent practical concerns have thus motivated the development of pricing rules that ensure lower levels of MWPs. Among them is the Join pricing rule (Ahunbay et al.,, 2022), which proposes a dual pricing problem that combines the objective of IP pricing (minimizing LLOCs) with minimizing MWPs. Specifically, for each participant it considers the maximum of LLOCs and MWPs in the objective, leading to a robust and computationally tractable pricing rule with adequate congestion signals and low MWPs simultaneously.
@@ 3 Data and Processing
This study examines welfare, prices, GLOCs, LLOCs, and MWPs for multiple zonal/nodal configurations, allocation rules, and pricing mechanisms. As discussed, we leverage the data released in the context of the European bidding zone review (ENTSO-E,, 2022; ACER, 2022b, ). We focus on Germany as arguably one of the most important bidding zones regarding clearing volume and its central position in the European grid. In this section, we provide an overview of the data set and our processing of it.
@@ 3.1 Bidding Zone Review Data
Unfortunately, some parts of the data were aggregated before their publication (e.g., local demand to country-wide demand), and some data were not revealed (e.g., a mapping of generators in the input files to their location in the grid model). We elaborate on these issues in the following. Without access to these proprietary data, the outcome of the ENTSO-E study cannot be exactly reproduced. However, it still provides a foundation against which we can validate our own results.
@@
The grid model for the German transmission system includes information on the system topology, components, and operational states. Each substation comprises multiple voltage levels, each of which can comprise several topological nodes that are connected with grid components / transmission lines. We consider power transformers as regular transmission lines and neglect their phase-shifting capability. For the grid constraints of our DCOPF in (4), we derive the per-unit susceptance $B_{nm}$ for each line from the given actual susceptance and base admittance (Grainger & Stevenson,, 1994). We also induce transmission limits on each power line, which we infer as the minimum of each line’s angular stability limit (inferred from its surge impedance loading and the St. Clair curve (St. Clair,, 1953)), voltage drop limit (inferred from voltage and reactance (Hao & Xu,, 2008)), and thermal limit (inferred from voltage and current on a three-phase system (Moreira et al.,, 2006)). We assume these transmission limits to be constant over time. Missing data, e.g., on the length and susceptance of lines, were replaced by the data in the JAO Static Grid Model where possible, and else inferred from the mean of all other lines.
The grid model consists of 2998 topological nodes. A significant part of these nodes represent breakers and disconnectors, and thus control devices that are neither associated with supply or demand nor changing voltage levels. Therefore, we disregard breakers and disconnectors and merge all topological nodes and their associated equipment at a given voltage level. Nodes not connected to the rest of the grid are merged with nodes associated with the same substation. After preprocessing, our grid model consists of 1687 nodes and 3232 lines (Schmitt,, 2023). In order to illustrate our results, we manually derived geographical coordinates for each substation from OpenStreetMap.
@@
For the demand side, the input files provide hourly aggregated load profiles for the entire country of Germany. Neither the nodal distribution of demand nor demand valuations are provided. Demand-side response was considered in the ENTSO-E study, but no related data was released. Consequently, we distribute demand proportional to each consumer’s base load that is available as part of the grid model. We thereby assume that demand fluctuations over time occur uniformly across all consumers (Schmitt,, 2023). We further assume that demand is exclusively price-inelastic.
@@
An additional challenge is the mapping of generators to corresponding network nodes. The grid data specify the broad category of generating units at each node (e.g., thermal, hydro, external injections) and their nominal capacities. While renewable resources could be mapped to the seller data by matching nominal capacities, the grid data lacks additional information regarding thermal plant types. We developed a heuristic to map thermal generators in the grid to their operational characteristics (such as minimum uptime constraints) (Schmitt,, 2023). Firstly, we match identifiers with the Kraftwerksliste (power plant list) provided by the Bundesnetzagentur to obtain each generator’s broad thermal generation type (e.g., lignite, gas, etc.). As nominal capacities do not match between grid and seller data, and seller data have a higher granularity (e.g., Gas CCGT, Gas OCGT, etc.), we solve an integer program to assign each entity in the grid model to exactly one corresponding seller of the same plant type, with the objective to minimize the aggregated differences in capacities and number of units per plant type. We provide this integer problem and the outcome of the generation matching in A While this assignment might not be exact, we argue that smaller mistakes in the assignment of generators to nodes have little impact on the more general questions raised in our paper that address the benefits of different zonal and nodal configurations. We provide our code files and a data description in an online appendix. The data are accessible through the ENTSO-E website.
@@ 3.2 Experimental Design
We implemented the models in Python 3.8 using Gurobi to solve optimization problems. We conducted experiments for eight weeks in 2009, as suggested by the ENTSO-E report (ENTSO-E,, 2022), for the German bidding zone. To maintain computational feasibility of computing hourly allocations and prices, we permitted a MIP gap of 5% for mixed-integer programs. Table 1 summarizes our experimental design and the scope of our analysis.
We first employ different allocation rules in accordance with the ENTSO-E report. These configurations range from a national single-zone model to a fully nodal model, as well as the ACER proposals for two to four bidding zones ACER, 2022b . Each allocation is associated with generation costs as well as redispatch costs to obtain a physically feasible outcome. After computing an allocation, different pricing rules, such as IP, CH, or Join pricing, can be applied. Each pricing rule implies different price levels as well as GLOCs, LLOCs, and MWPs for market participants. Prices were capped at EUR 100 per MWh to limit the impact of outliers. We also consider the Euphemia allocation and prices as the current implementation in European markets.
@@ 04 National IP Generation Costs
15 3 Zones GLOCs
16 4 Zones LLOCs
21 Nodal MWPs
@@ 4 Results
In this section, we summarize the main results of our analysis. We start with comparing average generation and total costs, before discussing price levels, GLOCs, LLOCs, and MWPs. We report aggregate statistics for all weeks under consideration and further pick the days of February 18, March 11, and November 23 for an in-depth analysis of representative days with low, medium, and high prices, respectively.
@@ 4.1 Generation and Redispatch Costs
Generation 33,507.00 33,507.68 33,507.37 33,516.13 33,510.40 36,008.27
Min-Cost Redispatch 5,124.61 5,120.69 5,127.95 5,117.03 5,145.39 0.00
Minimum Total Costs A 38,631.61 38,628.37 38,635.32 38,633.16 38,655.79 36,008.27
Min-Volume Redispatch 8,195.76 8,175.75 8,178.12 8,204.01 8,175.20 0.00
Minimum Total Costs B 41,702.76 41,683.43 41,685.49 41,720.14 41,685.60 36,008.27
@@
The numbers reported in Table 2 have realistic orders of magnitude as compared to those provided by the German Bundesnetzagentur (Bundesnetzagentur,, 2024). However, our goal is not to come up with a precise prediction of costs for 2025, but to obtain a relative comparison between different zonal configurations and nodal pricing. Adding network constraints cannot lead to lower objective function values in the cost minimization. As the national allocation rule neglects all transmission constraints, their generation costs are a lower bound for all other allocation rules. On average, the suggested zonal configurations require only slightly more generation costs, suggesting that adding few transmission constraints to the computation has little impact on the cost. In contrast, the nodal allocation rule includes the entire transmission network. Obviously, higher generation costs are required to reach an allocation with feasible power flows. On average, generation costs increase by 7.46% after including all network constraints for the nodal model compared to the national configuration before redispatch. After redispatch, the national configuration with minimum cost redispatch has 7.29% higher costs than the nodal model, and 15.8% higher costs with minimum volume redispatch. Note that we assume that all generators can be redispatched in our model, such that redispatch costs in reality might be higher. In comparison, the average daily redispatch costs in Germany between 2020 and 2022 amount to roughly kEUR 7,200 (Bundesnetzagentur,, 2022). The average daily redispatch costs in Germany in 2022 were even at around kEUR 11,600, having increased by almost 100% compared to 2021. We assume zero redispatch for a nodal system that satisfies $\Psi^{DC}$, although in practice there will always be some minimal level of redispatch due to changing weather conditions and the simplified representation of network flows in the DCOPF.

### v4

@@ 2.2.1 Zonal Market Clearing
The key idea of zonal pricing is to aggregate nodes into zones and only consider transmission flows across zones. This simplifies the constraint set $\Psi$ significantly, enhances computational scalability, and allows for portfolio bidding of larger market participants. An important parameter of a zonal power system is the determination of the capacities of interconnectors between the zones.
In the European market, two approaches have been implemented: the net transfer capacity (NTC) model and the flow-based market coupling (FBMC) methodology (NEMO Committee,, 2019). FBMC was introduced in Central Western Europe in 2015 and has since been extended to neighboring markets. In FMBC, the net export possible from one zone to another is limited by the remaining available margin (RAM) on the critical network elements connecting the zones. The RAM is determined by the maximum allowable flow (e.g., thermal capacity) subtracting the physical flow before the optimization (the base case), and a flow reliability margin to cover uncertainty in the capacity and allocation computations (Marien et al.,, 2013; Weinhold & Mieth,, 2024). The Flow-Based Market Coupling with Generation Shift Keys (FBMC-GSK) is the specific methodology currently being used in the EU. The way how the interconnection capacity between two price zones is calculated and allocated in FBMC-GSK depends on how various parameters are set and the critical network elements considered by the TSOs (Aravena et al.,, 2021), and cannot be predicted precisely for an eventual split of the German price zone. Furthermore, the FBMC approach necessitates access to a PTDF matrix that indicates how zone net position changes affect cross-zonal flows. Such a PTDF matrix is not included in the BZR data.
In our zonal market clearing implementation, we consider all interconnectors as critical network elements and model their physical characteristics via the DCOPF constraints (4). In line with earlier literature (Tiedemann et al.,, 2024; Wyrwoll et al.,, 2018; Voswinkel et al.,, 2019), we compute a base case that only takes 80% of the capacities of all interconnectors into account.Similar to the nodal model, this should take into account the $(N-1)$ security criterion (Christie et al.,, 2000). The set of feasible zonal solutions can be written as:
$\Psi^{Zonal}=\{(x,y):\sum_{s\in S}y_{s,zt}-\sum_{b\in B}x_{b,zt}=\sum_{n\in Z,m\notin Z}-B_{nm}(\theta_{nt}-\theta_{mt})\;\forall z\in Z,t\in T\}$ (5)
@@ 2.2.2 Redispatch
We simulate a cost-based redispatch. As redispatch should be cost-minimal, the approach is to find a modified allocation that satisfies the constraint set of $\Psi^{DC}$ and, at the same time, minimizes the required compensations to sellers for their changes in generation. Demand is excluded from redispatch, as it is current practice in Germany. More formally, with $(x^{Zonal},y^{Zonal})$ as the outcome of the zonal allocation and $f_{nmt}=B_{nm}(\theta_{nt}-\theta_{mt})$ obtained from (5) such that nodes $n$ and $m$ belong to two different zones, we solve the following problem:
@@
subject to $B_{nm}(\theta_{nt}-\theta_{mt})\leq f_{nmt}\ \ \forall n,m\in N,zone(n)\neq zone(m),t\in T$
$(x,y)\in\Psi^{DC}$
Here, $zone(n)$ denotes the zone to which node $n$ belongs. The objective of the optimization problem described in (6) is to adjust the zonal allocation $(x^{Zonal},y^{Zonal})$ in a manner that yields a DCOPF-feasible solution $(x,y)$ at minimal costs. This adjustment involves compensating sellers for any changes in their dispatched quantities during redispatch. For each seller $s$, this compensation is determined by the difference between the costs they incur under the initial zonal allocation $c_{s}(y_{s}^{Zonal})$ and the costs under the final DCOPF-feasible allocation $c_{s}(y_{s})$. This assumes that every generator can be redispatched, which might not be the case in practice. Note that the constraints on cross-zonal line capacities from the zonal model are also taken into account in the redispatch computation. The objective of (6) is to minimize the total costs associated with these adjustments. As part of our experiments we also report results for redispatch that minimizes volumes instead of costs, i.e., the redispatch objective function changes to $\min_{x,y}\;\sum_{s\in S}\lvert y_{s}-y_{s}^{Zonal}\rvert$.
@@ 2.3 Pricing on Non-Convex Markets
Consequently, market operators have to resort to heuristic pricing rules that aim at approximating Walrasian equilibrium prices. Many such pricing rules have been proposed (Liberopoulos & Andrianesis,, 2016), each compromising on the properties of Walrasian equilibria differently. In this work, we focus on Convex Hull (CH) pricing (Hogan & Ring,, 2003; Gribik et al.,, 2007) and Integer Programming (IP) pricing (O’Neill et al.,, 2005) as two established pricing rules in practice, as well as on the Join pricing rule (Ahunbay et al.,, 2024) as a novel multi-objective approach to pricing. We also consider the Euphemia pricing rule as it is applied in European electricity markets.
@@ 2.3.2 IP Pricing
$\text{LLOC}_{s}$ $=\hat{u}^{\prime}_{s}(p)-u_{s}(y|p)\text{ with }\hat{u}^{\prime}_{s}(p)=\hat{u}_{s}(p)\text{ s.t. }u=u^{*}$
By convexity of $v_{b}(x_{b})$, a buyer’s LLOCs exactly equal their GLOCs. Besides computational tractability, IP prices accurately reflect the marginal value of transmission capacity (Yang et al.,, 2019). Specifically, price differences among nodes arise solely when the network experiences congestion. It has been shown that this property immediately corresponds minimizing LLOCs (Ahunbay et al.,, 2024), making them a significant indicator of good congestion signals.
@@ 2.3.3 Join Pricing
$\text{MWP}_{b}$ $=\max\{-u_{b}(x|p),0\}$
$\text{MWP}_{s}$ $=\max\{-u_{s}(y|p),0\}$
Bids that require MWPs are also known as paradoxically accepted bids. MWPs can be regarded as another subset of GLOCs, referring only to participants’ incentives to not participate in the market at all. However, both CH and IP pricing can experience high levels of MWPs. Recent practical concerns have thus motivated the development of pricing rules that ensure lower levels of MWPs. Among them is the Join pricing rule (Ahunbay et al.,, 2024), which proposes a dual pricing problem that combines the objective of IP pricing (minimizing LLOCs) with minimizing MWPs. Specifically, for each participant it considers the maximum of LLOCs and MWPs in the objective, leading to a robust and computationally tractable pricing rule with adequate congestion signals and low MWPs simultaneously.
@@ 3 Data and Processing
This study examines welfare, prices, and MWPs for multiple zonal/nodal configurations, allocation rules, and pricing mechanisms. As discussed, we leverage the data released in the context of the European bidding zone review (ENTSO-E,, 2022; ACER, 2022b, ). We focus on Germany as arguably one of the most important bidding zones regarding clearing volume and its central position in the European grid. In this section, we provide an overview of the data set and our processing of it.
@@ 3.1 Bidding Zone Review Data
Unfortunately, some parts of the data were aggregated before their publication (e.g., local demand to country-wide demand), and some data were not revealed (e.g., a mapping of generators in the input files to their location in the grid model). Without access to these proprietary data, the outcome of the ENTSO-E study cannot be exactly reproduced. However, we could assign most of generators to locations via OpenStreetMap. [^8] https://www.openstreetmap.de/
Let us discuss the data in more depth. The topology of the grid, defined by its nodes and connections, dictates the path electricity takes from generation to consumption, with substations facilitating the necessary voltage transformations along the way. Substations act as critical nodes in the network that manage voltage levels for efficient power flow. A substation might handle multiple voltage levels. For example, a transmission substation might receive power at a very high voltage (e.g., 220kV or higher) and step it down to a lower high voltage (e.g., 110kV) for further transmission or to a medium voltage (e.g., 10-35kV) for distribution to urban or industrial areas.
The ENTSO-E dataset provides a model of Germany’s power grid, detailing 834 substations. Each of these substations encompasses one or more voltage levels, summing up to 1697 voltage levels at substations in the dataset. These voltage levels cover the spectrum from transmission ($\geq$ 220kV) to distribution ($\leq$ 110kV) and are composed of several topological nodes. Altogether, there are 2898 topological nodes listed in the data sets. Additionally, the dataset includes details for 100 substations located in the neighboring countries, which have been integrated into our analysis. For our study, we manually determined the geographical locations of all substations and computed prices for all voltage levels on these substations. After some pre-processing, which involved removing disconnected nodes, breakers, and disconnectors, our German grid model consists of 1670 nodes and 2470 lines (Schmitt,, 2023). Technically, supply and demand on different voltage levels of substations do not necessarily need to be the same. However, it turned out that prices computed for voltage levels on a substation were identical or very close in the ENTSO-E dataset, which is also the case for the nodal prices provided by ENTSO-E. Overall, the average prices that we computed were not significantly different from those in the BZR dataset for substations. There were differences on individual nodes, which is probably due to the fact that the prices computed by ENTSO-E are based on the duals of a linear programming relaxation, while we use the mixed integer linear program and pricing rules as they are defined above.
@@
For the grid constraints of our DCOPF in (4), we derive the per-unit susceptance $B_{nm}$ for each line from the given actual susceptance and base admittance (Grainger & Stevenson,, 1994). We also induce transmission limits on each power line, which we infer as the minimum of each line’s angular stability limit (inferred from its surge impedance loading and the St. Clair curve (St. Clair,, 1953)), voltage drop limit (inferred from voltage and reactance (Hao & Xu,, 2008)), and thermal limit (inferred from voltage and current on a three-phase system (Moreira et al.,, 2006)). We assume these transmission limits to be constant over time. Missing data, e.g., on the length and susceptance of lines, were replaced by the data in the JAO Static Grid Model where possible, and else inferred from the mean of all other lines. As discussed earlier, we reserved a buffer of 20% on the transmission limits to account for secure operations of the electricity grid.
@@
For the demand side, the input files provide hourly aggregated load profiles for the entire country of Germany. Neither the nodal distribution of demand nor demand valuations are provided. Demand-side response was considered in the ENTSO-E study, but no related data was released. We distributed demand proportional to each consumer’s base load that is available as part of the grid model. We thereby assume that demand fluctuations over time occur uniformly across all consumers (Schmitt,, 2023). We further assume that demand is price-inelastic, as was done in the study by ENTSO-E.
@@
An additional challenge is the mapping of generators to corresponding network nodes. The grid data specify the broad category of generating units at each node (e.g., thermal, hydro, external injections) and their nominal capacities. While renewable resources could be mapped to the seller data by matching nominal capacities, the grid data lacks additional information regarding thermal plant types. We assigned thermal generators in the grid to their operational characteristics (such as minimum uptime constraints) (Schmitt,, 2023). Firstly, we match identifiers with the Kraftwerksliste (power plant list) provided by the Bundesnetzagentur to obtain each generator’s broad thermal generation type (e.g., lignite, gas, etc.). As nominal capacities do not match between grid and seller data, and seller data have a higher granularity (e.g., Gas CCGT, Gas OCGT, etc.), we solve an integer program to assign each entity in the grid model to exactly one corresponding seller of the same plant type, with the objective to minimize the aggregated differences in capacities and number of units per plant type. We provide this integer program and the outcome of the generation matching in A. While this assignment might not be exact, we argue that smaller mistakes in the assignment of generators to nodes have little impact on the more general questions raised in our paper that address the benefits of different zonal and nodal configurations. The data are accessible through the ENTSO-E website.
@@ 3.2 Experimental Design
We implemented the models in Python 3.8 using Gurobi to solve optimization problems. We conducted experiments for eight weeks in 2009, as suggested by the ENTSO-E report (ENTSO-E,, 2022), for the German bidding zone. To maintain computational feasibility of calculating hourly allocations and prices, we permitted a MIP gap of 5% for mixed-integer programs. Table 1 summarizes our experimental design and the scope of our analysis.
We first employ different allocation rules in accordance with the ENTSO-E report. These configurations range from a national single-zone model to a fully nodal model, as well as the ACER proposals for two to four bidding zones ACER, 2022b . Each allocation is associated with generation costs as well as redispatch costs to obtain a physically feasible outcome. After computing an allocation, different pricing rules, such as IP, CH, or Join pricing, can be applied. Each pricing rule implies different price levels as well as MWPs for market participants. Prices were capped at 100 EUR/MWh to limit the impact of outliers. Note that prices higher than 100 EUR/MWh are regarded as outliers also in the BZR Report (ENTSO-E,, 2022). We also consider the Euphemia allocation and prices as the current implementation in European markets.
@@ 04 National IP Generation Costs
15 3 Zones MWPs
16 4 Zones
21 Nodal
@@ 4 Results
In this section, we summarize the main results of our analysis. We start with comparing average generation and total costs, before discussing price levels and MWPs. We report aggregate statistics for all weeks under consideration and further pick the days of February 18, March 10, and November 23 for an in-depth analysis of representative days with low, medium, and high prices, respectively.
@@ 4.1 Generation and Redispatch Costs
Generation 33,507.00 33,702.87 33,697.10 33,578.77 33,717.41 36,150.7
Min-Cost Redispatch 3,567.84 3,421.81 3,457.25 3,546.4 3,449.39 0
Minimum Total Costs A 37,074.84 37,124.68 37,154.35 37,125.17 37,166.8 36,150.7
Min-Volume Redispatch 5,736.95 5,506.98 5,598.22 5,694.76 5,560.43 0
Minimum Total Costs B 39,243.95 39,209.85 39,295.32 39,273.53 39,277.84 36,150.7
@@
The numbers reported in Table 2 have realistic orders of magnitude as compared to those provided by the German Bundesnetzagentur (Bundesnetzagentur,, 2024). However, our goal is not to come up with a precise prediction of costs for 2025, but to obtain a relative comparison between different zonal configurations and nodal pricing. Adding network constraints cannot lead to lower objective function values in the cost minimization. As the national allocation rule neglects all transmission constraints, their generation costs are a lower bound for all other allocation rules. On average, the suggested zonal configurations require only slightly more generation costs, suggesting that adding few transmission constraints to the computation has little impact on the costs. In contrast, the nodal allocation rule includes the entire transmission network. Obviously, higher generation costs are required to reach an allocation with feasible power flows. On average, generation costs increase by 7.46% after including all network constraints for the nodal model compared to the national configuration before redispatch.
After redispatch, the national configuration with minimum cost redispatch has 2.5% higher costs than the nodal model, and 13.65% higher costs with minimum volume redispatch. Note that we assume that all generators can be redispatched in our model, such that redispatch costs in reality might be higher. In comparison, the average daily redispatch costs in Germany between 2020 and 2022 amount to roughly kEUR 7,200 (Bundesnetzagentur,, 2022). The average daily redispatch costs in Germany in 2022 were even at around kEUR 11,600, having increased by almost 100% compared to 2021. We assume zero redispatch for a nodal system that satisfies $\Psi^{DC}$, although in practice there will always be some minimal level of redispatch due to changing weather conditions and the simplified representation of network flows in the DCOPF. At first sight, it might be curious that redispatch costs can even increase with additional zones. B provides a simplified example to explain how this can happen.